OpenStudy (anonymous):

A student studying for a vocabulary test knows the meanings of 10 words from a list of 24 words. If the test contains 10 words from the study list, what is the probability that at least 8 of the words on the test are words that the student knows? (Round your answer to three decimal places.)

OpenStudy (kropot72):

\[P(8\ known)=\frac{10C8\times14C2}{24C10}\] \[P(9\ known)=\frac{10C9\times14}{24C10}\] \[P(10\ known)=\frac{1}{24C10}\] After calculating the above 3 values of probability, you need to add them to find the probability that at least 8 of the words on the test are words that the student knows.